A dog's fur is combed, and the comb gains a charge of 8.0 nC. Find the electric force between the fur and comb when they are 2.0 cm apart. 



1
Expert's answer
2021-09-13T17:52:50-0400

There is no exact answer to this question. Since both dog and comb can not be considered as a point charges if they only 2cm apart.

But, we can attempt to apply Coulomb's law anyway, keeping in mind that the result will be definitely wrong.


"F = k\dfrac{q_1q_2}{r^2}"

where "q_1 = q_2 = 8\times 10^{-9}C" are the modulus of the charges on comb and fur (they are equal according to the charge conservation law), "r = 2cm = 0.02m" is the distance, and "k = 9\times 10^9Nm^2/C^2" is the constant. Thus, obtain:

"F = 9\times 10^9 \cdot \dfrac{(8\times10^{-9})^2}{(0.02)^2} \approx 1.4\times 10^{-3}N"

Answer. "1.4\times 10^{-3}N" (taking into account the initial note).